A cream pie (m = 0.80 kg) is fired out of a circus cannon with an initial speed of 19 m/s from a height of 2.50 m. The pie follows a parabolic path, eventually hitting a clown in the knees with a speed of 20 m/s during its descent.

a) Calculate the initial kinetic energy of the cream pie as it leaves the cannon.

b) Calculate the initial gravitational potential energy of the cream pie as it leaves the cannon.

c) Calculate the final kinetic energy of the cream pie just as it hits the clown’s knees.

d) Use the Law of Conservation of Energy to calculate the final gravitational potential energy of

the cream pie when it hits the clown’s knees.

e) Calculate the final height of the cream pie as it hits the clown’s knees. 


1
Expert's answer
2021-01-25T14:05:18-0500

a)


"K_i=0.5(0.8)(19)^2=144.4\ J"

b)


"P_i=(0.8)(9.8)(2.5)=19.6\ J"

c)


"K_f=0.5(0.8)(20)^2=160\ J"

d)


"P_f=144.4+19.6-160=4\ J"

e)


"(0.8)(9.8)h=4\\h=0.51\ m"